OpenStudy (anonymous):

a company invests $15,000 in an account that compounds interest annually. after two years the account is worth $16,099.44. us the function r is the annual interest rate, p is the pricipal and a is the amount of money after t years. what is the interest rate on the account? A=P(1+r)t

OpenStudy (mathmate):

A=P(1+r)^t For t=2, 1+r=sqrt(A/P) solve for r.